Twin brothers, Billy and Bobby, can mow their grandparents lawn together in 89 minutes. Billy could mow the lawn in 30 minutes less time than it would take Bobby. How long would it take Bobby to mow by himself?
Answer by ad_alta(240)

Let Billy's rate be 'i' and Bobby's rate be 'o.' Then 1/i=1/o-30. Also, (i+o)*89=1. The first equation tells us that i=-o/(30i+1). So (o-o/(30*o-1))*89=1. That means o=0.005148. So it would take him 1/o or 194.25 minutes.